OpenStudy (anonymous):

\[\sqrt{x^2 + 4}\] Find the domain and range

OpenStudy (apoorvk):

Okay so, I feel... anyways... That thing under the sqrt, 'x+4' has to equal or greater than '0' right??

OpenStudy (anonymous):

sorry there is a change in question

OpenStudy (anonymous):

\[\sqrt{x^2 + 4}\]

OpenStudy (apoorvk):

yeah that as well. x^2+4 should be greater than or equal to zero. so can you set up an inequality for this?

OpenStudy (anonymous):

x^2 + 4 >= 0 x^2 > = -4

OpenStudy (anonymous):

Domain = R becoz we can put any numbers

OpenStudy (apoorvk):

Exactly!

OpenStudy (anonymous):

since there is sq term - ve will be + ve

OpenStudy (anonymous):

so wat abt Range??

OpenStudy (anonymous):

Range = [2 , infinity)

OpenStudy (apoorvk):

Now for the range - at x=0, is when the function attains its minimum value - what's that?

OpenStudy (anonymous):

is that correct

OpenStudy (apoorvk):

Yeah, you're right with the range too!

OpenStudy (anonymous):

thxx for helping!

OpenStudy (apoorvk):

No worries mate ;)